Please identify the type of conic section whose equation is given and find the vertices and foci.
y^2 - 8y = 6x - 16

Answer by stanbon(75887)

Please identify the type of conic section whose equation is given and find the vertices and foci.
y^2 - 8y = 6x - 16
---
y^2-8y+? = 6x-16 + ?
---
Complete the square on the y-terms.
---
y^2-8y+4^2 = 6x-16+4^2
---
Factor:
(y-4)^2 = 6(x-0)
----
4p = 6
p = 3/2
-----
Type: parabola opening to the right
Vertex: (4,0)
Focus: (4+3/2,0) or (5.5,0
==============================
Cheers,
Stan H.
=============